the result of a magazine survey showed that 42 out of 80 people plan to vote for Mr targeted for president.which is the best prediction of the total number of votes he will receive if 2000 people vote
Answer by stanbon(75887)

the result of a magazine survey showed that 42 out of 80 people plan to vote for Mr targeted for president.which is the best prediction of the total number of votes he will receive if 2000 people vote
-----------
Use a proportion::
x/2000 = 42/80
-----
x = 2000(21/40)
-----
x = 50*21
---
x = 1050
----------------
Cheers,
Stan H.
-----------------